OpenStudy (anonymous):

A sporting goods store is having a 15% off sale on all items. Which functions can be used to find the sale price of an item that has an original price of x ? You can choose more than one ƒ(x) = x - .15x Sale = Original - 15 ƒ(x) = 1.15x Sale = Original - .15(Original) y = .85x pls help pt 2

OpenStudy (mathstudent55):

The original price is x. The discount is 15%. How do you write 15% of x as an expression?

OpenStudy (mathstudent55):

In math, "of" usually means "times". 15% of x means 15% * x. 15% can be written as a decimal as 0.15 That means that 15% * x = 0.15x Do you understand so far?

OpenStudy (anonymous):

yeah kinda

OpenStudy (mathstudent55):

Here's an example. Maybe with an example it will become clearer. Let's say one item costs $100. 15% of $100 is 15% * 100 = 0.15 * 100 = 15 15% of 100 is 15. Ok?

OpenStudy (mathstudent55):

Ok. Let's continue. If you multiply a number by 0.15, you are finding 15% of the number. If the price is x, then 0.15x is 15% of x. The discount is 15%, so once you find the discount, you must subtract the discount from the original price to find the discounted price. The original price is x. The discount is 0.15x The final price is x - 0.15x As a function, you can write: f(x) = x - 0.15x This is an answer Since x and 0.15x are like terms, they can be combined together. f(x) = 1x - 0.15x f(x) = 0.85x This is also an answer With this we conclude that the sale price is the original price minus 0.15 of the original price subtracted from it. sale price = original price - 0.15(original price) This is also an answer